OpenStudy (anonymous):

How do you factor 6v^2 + 48v + 72?

OpenStudy (unklerhaukus):

first; divide by six

OpenStudy (anonymous):

48 divided by 6 is 8 and 72 divided by 6 is 12 and 6 divided by 6 is 1

OpenStudy (anonymous):

This is assuming of course that the equation above equals zero, or else you will have to divide both sides (essentially, you are anyways). So now the equation is \[v^2+8v+12=0\] Do you know how to factor this?

OpenStudy (anonymous):

is it (v + 6) (v + 2) ?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

what about x^2 - 4x - 12?

OpenStudy (unklerhaukus):

so \[ 6v^2 + 48v + 72=6(v+6)(v+2)\]

OpenStudy (anonymous):

ya that's what I got

OpenStudy (unklerhaukus):

for \[x^2 - 4x - 12\] the difference between two numbers is 4 and their product is -12

OpenStudy (anonymous):

Same way: Two numbers that multiply to equation -12, and add up to -4. In this case, -6 and 2 \[(x-6)(x+2)\]

OpenStudy (anonymous):

um 5w^2 - 20?

OpenStudy (anonymous):

I don't know how to factor this without having a middle number

OpenStudy (anonymous):

Factor out the 5 first.

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

for \[5w^2 - 20\] first factor out the multiple of five

OpenStudy (anonymous):

\[5(w^2-4)\] That is the pattern known as the difference of two squares. w is a square, 4 is a square. It follows the pattern: \[(a^2-b^2)=(a+b)(a-b)\] Apply that two \(w^2-4\) and see what you get.